HOMEWORK 1. If the capacitor is initialy uncharged, calculate the time atfher the switch is closed mecessary for it to acquire 632% of its final charge in this circuit? Va-30 volts R-500 kΩ

Help please! and workout problems in detail!!
Answer #1

given,

V = 30 V

R = 500 * 10^3 ohm

C = 35 uF

Vc = Vs * (1 - e^(-t/RC))

0.632 = (1 - e^(-t/(500 * 10^3 * 35 * 10^-6)))

time = 17.49 sec